OpenStudy (anonymous):

In a fourth year of a school,there were 156 students who took both mathematics and Science.They all passed at least one subjectand 75 passed both subjects. If twice as many passed Science as passed Mathematics, find out how many passed Mathematics only.

OpenStudy (anonymous):

Do you know how to start it off?

OpenStudy (gorv):

ok it is venne diagram Q

OpenStudy (gorv):



OpenStudy (gorv):

let only passed in math=x and passed in only science=y

OpenStudy (anonymous):

no

OpenStudy (gorv):

x=2y

OpenStudy (anonymous):

Will I be correct if I say y=2x/

OpenStudy (gorv):

75 student passed both math and science remaining student passed in either math or science=156-75=81

OpenStudy (gorv):

yeah y=2x u r right x+y=81 and y=2x x+2x=81 3x=81 x=81/3 x=27 y=54

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (gorv):

welcome